P(2,3) L: Kx-2y+k=0
Found 2 solutions by ikleyn, tommyt3rd:
Answer by ikleyn(52798)

.

Substitute the coordinates   x= 2,   y= 3   into the equation and find   k.



Answer by tommyt3rd(5050)

Unless we assume K=k there is no unique solution to this problem. So let's get started...
(2,3) implies x=2, y=3
Using substitution we can say kx-2y+k=0 becomes k(2)-2(3)+k=0. Then
2k-6+k=0
2k-6+6+k=0+6
3k=6
3k/3=6/3
k=2
which leads to L: 2x-2y+2=0
and we can simplify to get L: x-y+1=0